OpenStudy (anonymous):

if a line contains a point (-3,5) and has slope m=-2/3, then the equation of the line in standard form is

OpenStudy (amity):

You can apply point slope form here as y-y1=m(x-x1). y-(-3)=-2/3(x-5),,,,,3(y+3)=-2(x-5) ,3y+9=10-2x,,,3y+2x+9-10=0,2x +3y-1=0 is general equation of straight line

OpenStudy (amity):

You can apply point slope form here as y-y1=m(x-x1). y-(-3)=-2/3(x-5),,,,,3(y+3)=-2(x-5) ,3y+9=10-2x,,,3y+2x+9-10=0,2x +3y-1=0 is general equation of straight line

OpenStudy (anonymous):

thank you so much! i appreciate it <3

OpenStudy (amity):

Welcome loveangel

OpenStudy (anonymous):

the pressure P (in punds per square foot) at adepth of d (in feet) below the surface of the water is given by the function p(d)=63.9d what is the pressure at a depth of 100 feet?

OpenStudy (anonymous):

that's the last question i have if you don't mind helping me with it please.. i tried to solve but i'm not coming with the right answer!